OpenStudy (anonymous):

find where the average rate of change is the same as the instantaneous rate of change for f(x)=x^3+x=1 over [0,2]

OpenStudy (anonymous):

Use LMVT(Lagrange Mean Value theorem) f'(c)=((f(b)-f(a))/(b-a) 3c+1=(11-(-1))/(2-0) 3c+1=6 c=5/3

OpenStudy (anonymous):

So that's all I have to do?

OpenStudy (anonymous):

Yes